OpenStudy (anonymous):

Can someone please help me? I'm horrible at math ! 2x + y = 11 3y - 4y = 11 you have to solve the problem using substitution. please help.

OpenStudy (anonymous):

3y - 4y= -y

OpenStudy (anonymous):

canu tell me wat is -4y+3y

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

Solve 2x + y = 11 for y. y = 11 - 2 x Replace each y in the second equation with 11 - 2 x and solve the result for x. Can you take it from here ?

OpenStudy (anonymous):

it would be 33/7

OpenStudy (wolf1728):

y = 11-2x 3y - 4y = 11 3•(11-2x) -4•(11-2x) = 11 Does that help?

OpenStudy (wolf1728):

33 -6x -44 +8x = 11 -11 + 2x = 11 2x = 22 x = 11 y=-11